OpenStudy (anonymous):

Two bowling balls are sitting on the ground. The Blue ball has a mass of 5kg. The Orange ball has a mass of 10kg. Both balls are rolled (pushed) with a force of 10N. Determine the acceleration of each ball. Explain why the Orange ball will have a lower acceleration than the Blue ball. a. Acceleration of the Blue ball b.Acceleration of the Orange ball c.explain why the orange ball has a lower acceleration

OpenStudy (anonymous):

f=m*a for orange ball f/m=a 10N/10kg=1m/s sq. for blue ball 10N/5kg=2m/s sq

OpenStudy (anonymous):

see orange ball has more mass than blue ball therefore according to the law of inertia which says more the mass of object, more the tendency to stay in motion or rest hence f/m=a so more the mass less the accelaration

OpenStudy (anonymous):

thanks so much